# SciReplicate-Bench: Benchmarking LLMs in Agent-driven Algorithmic Reproduction from Research Papers

### v1

@@ 4.1 Results on SciReplicate-Bench
Table 3 displays Sci-Reproducer’s evaluation results and contributions of code/paper agent. The “No Agent”directly prompts the LLM to generate code based solely on the algorithm LaTeX code and function signature. “No Paper Agent” allows the LLM to use code agent actions, i.e., website and repository research and compiler incorporation, but restricts access to paper agent actions. “No Code Agent” grants access to paper agent actions but blocks Code Agent capabilities. The results offer key insights, discussed in the following.
LLMs struggles on SciReplicate-Bench Most LLMs perform poorly, achieving less than 0.1 execution accuracy without using the agent to examine literature and repository contexts. With enhancement of Sci-Reproducer, these LLMs showed notable improvements, with an average increase of 0.181 in execution ACC and 0.057 in CodeBLEU, although even the best-performing model, Claude-Sonnet-3.7, only achieved 0.390 execution accuracy. This highlights the exceptional challenge presented by our SciReplicate-Bench.
LLMs can comprehend algorithm logic Most models are capable of understanding the core implementation logic of target algorithms (indicated by reasoning graph accuracy), even without any external assistance, with an average score of 0.731 in the “No Agent” setting. Both the Paper Agent and Code Agent further enhance the algorithm understanding, i,e., leading to an average increase of 0.013 and 0.049, respectively; using both agents together results in an improvement of 0.060. These can be explained that Paper Agent’s retrieved contextual information from the surrounding literature can help the model comprehend the theoretical and methodological aspects; the Code Agent extracts relevant code snippets and dependencies from code repositories for the implementation logic understanding.
@@
Overthinking leads to limited improvement in reasoning LLMs The performance gains of reasoning LLMs with Sci-Reproducer are relatively limited. Specifically, they show an average improvement of 0.13 in execution accuracy, compared to a higher average improvement of 0.212 observed in non-reasoning ones. For recall metrics, reasoning LLMs achieve average gains of 0.243, 0.061, and 0.041, respectively. In contrast, non-reasoning LLMs demonstrate notably higher improvements of 0.560, 0.345, and 0.135 on the same metrics. This discrepancy implies that reasoning LLMs tend to rely heavily on their internal reasoning capabilities, rather than pre-defined actions. This tendency highlights a limitation refer to as “overthinking” (Cuadron et al., 2025; Sui et al., 2025), which we further analyze in the next subsection.
@@ 4.2 Tool Usage Analysis
Figure 2 presents the number of times each LLM invokes actions on the full dataset with Sci-Reproducer. The first four actions are code-related actions, while the last three are paper-related actions. We observe the following:
  • Regarding code-related actions, reasoning LLMs use “SearchFile”, “SearchCodeItem”, and “SearchWeb” an average of 25.0, 3.3, and 0.0 times, respectively. In comparison, non-reasoning LLMs use these actions significantly more often, with averages of 210.4, 68.2, and 16.8 times, respectively. This suggests that reasoning models tend to rely more on internal deliberation rather than retrieving external information. In contrast, the invocation frequency of “Compiler” is notably higher for reasoning models, indicating that they require more attempts to correct syntactic errors. Such over-reliance on internal reasoning hurts their overall performance: the execution accuracy of models such as o3-mini-high and o3-mini-low is comparable to that of gpt-4o-mini, despite their theoretical advantages.
  • A similar trend is observed for paper-related actions. Reasoning LLMs use “SearchPaper”, “SearchSection”, and “SearchLiterature” an average of 56.3, 70.0, and 20.0 times, respectively. In contrast, non-reasoning LLMs make significantly greater use of these actions, with average usage counts of 244.8, 188.4, and 58.0, respectively. Furthermore, LLMs are more inclined to extract information from the target paper rather than from related literature. On average, all LLMs invoke the actions “SearchPaper” and “SearchSection”—which retrieve information from the target paper—174.1 and 144 times, respectively, while “SearchLiterature”, which accesses related literature, is used only 43.8 times on average.
@@ 4.3.1 Syntax Errors
Table A1 shows the syntax error rates for each model across different configurations. Without the Code Agent, syntax errors occurred at rates of 80.3% (“NoAgent”) and 83.3% (“Paper Agent”). After implementing the Code Agent, these error rates dropped significantly to 29.4% (“Code Agent”) and 24.9% (“Sci-Reproducer”). The remaining syntax errors mainly result from incorrectly using repository dependencies. This occurs because our approach, unlike human developers, cannot dynamically access variable values through an compiler during the code generation process.
@@ 4.3.2 Logic Errors
5 Related Work
Our work lies at the intersection of AI for scientific discovery and LLM-based code generation. While prior studies (Wang et al., 2023; Ghafarollahi & Buehler, 2024; Si et al., 2024) explore LLMs for hypothesis generation, their ability to generate code for hypothesis verification remains underexplored. Existing code generation benchmarks focus on text-to-code tasks and ML software engineering, but none evaluate LLMs’ ability to reproduce algorithm or model description from real-world scientific papers. Additional details are provided in Appendix A.
@@ 6 Conclusion
We evaluate LLMs’ ability to replicate algorithms described in recent NLP papers. To support this, we introduce SciReplicate-Bench, a benchmark with rich annotations, and Sci-Reproducer, a multi-agent framework for bridging algorithm understanding and code generation. We assess performance using reasoning graph accuracy and standard implementation metrics. Results show the task is highly challenging, with failures largely caused by missing or inconsistent algorithm descriptions.
@@ 7 Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2).
@@
Appendix A Extended Discussion of Related Work
A.1 AI for Automating Scientific Discovery
Numerous research efforts aim to accelerate scientific discovery using large language models (LLMs) (Jumper et al., 2021; Wang et al., 2023; Ghafarollahi & Buehler, 2024; Si et al., 2024; Schmidgall et al., 2025; Lu et al., 2024; Sethi et al., 2017). Some studies focus on scientific hypothesis generation: Wang et al. (2023); Ghafarollahi & Buehler (2024) explore the potential of LLMs to propose novel research concepts, while Si et al. (2024) suggests that AI-generated ideas can, in some cases, surpass those of human researchers in originality.
In parallel, Schmidgall et al. (2025); Lu et al. (2024) introduce AI systems capable of automating the end-to-end research workflow, encompassing idea generation, validation, and manuscript drafting. Although these systems leverage LLMs to implement algorithms, the algorithms themselves are synthesized by the models, rather than derived from peer-reviewed scientific publications. As a result, they lack the complexity and scientific rigor typically found in real research algorithms. Moreover, these works do not verify whether the generated code faithfully implements the underlying algorithmic logic.
In contrast, our work is the first to investigate the ability of LLMs to reproduce algorithms proposed in peer-reviewed academic papers, bridging the gap between natural language understanding and faithful scientific code generation.
A.2 LLMs for Code Generation
Code generation has been a prominent research topic in natural language processing (NLP), giving rise to a range of benchmarks and methodologies. Several works evaluate models’ ability to generate code from natural language descriptions (Chen et al., 2021; Jain et al., 2024; Austin et al., 2021; Hendrycks et al., 2021; Liu et al., 2022). Although state-of-the-art models have achieved high performance on many of these benchmarks, they still fall short of fully automating the role of a software engineer. To address realism and complexity, SWE-bench (Jimenez et al., 2023) introduces tasks based on actual pull requests from open-source repositories. However, most of these benchmarks are grounded in the domain of general software engineering.
Some recent efforts have shifted focus toward machine learning-specific software engineering benchmarks (Liu et al., 2023; Huang et al., 2023; Chan et al., 2024). Nevertheless, their goals often involve implementing algorithms proposed by the models themselves or solving relatively simple tasks, which do not require the depth of algorithmic understanding or rigorous paper analysis involved in reproducing algorithms from peer-reviewed publications.
LLMs have demonstrated strong capabilities in complex code generation, particularly when equipped with tool-use strategies. Methods such as ToolFormer (Schick et al., 2023), KwaiAgents (Pan et al., 2023), CodeAgent (Zhang et al., 2024a) and ToolCoder (Zhang et al., 2023a) showcase the potential of LLMs to solve challenging tasks through effective tool integration.
Despite these advances, no existing system is specifically designed for reproducing algorithms from academic papers. Current approaches are not directly applicable in this setting, as they lack dedicated components for comprehending, interpreting, and aligning scientific descriptions with executable code.
Appendix B Details of the Annotation Process
@@
We curated papers from top-tier NLP conferences in 2024, including ACL, EMNLP, and COLING. Using a web crawler, we collected accepted paper titles and employed the PapersWithCode API [^4] https://paperswithcode.com/api/v1/docs/ to identify those with open-source implementations. For each identified paper, we retrieved corresponding GitHub repository links and metadata (e.g., stars, issues, release dates) via the GitHub REST API [^5] https://docs.github.com/en/rest?apiVersion=2022-11-28.
@@
In the final step, annotators created a verification suite with 10 test cases drawn from the original datasets used in each repository. Given the inherent randomness in many NLP implementations and potential machine-related variability, we addressed reproducibility from two angles:
@@
  • Controlling machine variability: users were instructed to run both reference and generated code locally to eliminate discrepancies caused by system-level differences.
@@
Appendix C Details of the Task Categories
@@
Appendix D Details of the Actions
@@
For each Python file in the code repository, we utilize the Python AST  [^7] https://docs.python.org/3/library/ast.html package to parse the file and extract all defined classes, functions, and global variables. Unlike embedding-based code search methods (Zhang et al., 2024c; 2023b), the Code Agent in our framework directly provides the name of a code item. The tool then returns the corresponding definition if it exists; otherwise, it returns an empty response.
@@ Appendix E Figures and Tables
Figure A3: An overview of the SciReplicate-Bench.

### v2

@@ 5.1 Results on SciReplicate-Bench
Table 3 displays Sci-Reproducer’s evaluation results and contributions of Code/Paper Agent. The “No Agent”directly prompts the LLM to generate code based solely on the algorithm description and function signature. “No Paper Agent” allows the LLM to use Code Agent actions, but restricts access to Paper Agent actions. “No Code Agent” grants access to Paper Agent actions but blocks Code Agent capabilities. The results offer key insights, discussed in the following.
LLMs struggles on SciReplicate-Bench Most LLMs perform poorly, achieving less than 0.1 execution accuracy without using the agent to examine literature and repository contexts. With enhancement of Sci-Reproducer, these LLMs show notable improvements, with an average increase of 0.181 in execution ACC and 0.057 in CodeBLEU, although even the best-performing model, Claude-Sonnet-3.7, only achieved 0.390 execution accuracy. This highlights the exceptional challenge presented by our SciReplicate-Bench.
LLMs can comprehend algorithm logic LLMs demonstrate strong algorithmic comprehension capabilities, as evidenced by reasoning graph accuracy scores averaging 0.716 even without agent assistance. The addition of individual agents provides modest but consistent improvements: the Paper Agent increases understanding by 0.009 on average, while the Code Agent contributes a larger gain of 0.036. Combined agent deployment yields a cumulative improvement of 0.037. These enhancements stem from complementary mechanisms: the Paper Agent strengthens theoretical comprehension by gathering relevant contextual information from academic literature, while the Code Agent facilitates practical understanding through extraction of pertinent code patterns and dependency structures from repositories.
@@
Overthinking leads to limited improvement in reasoning LLMs Reasoning LLMs exhibit more modest performance gains when using Sci-Reproducer. While they achieve an average execution accuracy improvement of 0.13, non-reasoning models demonstrate substantially larger gains of 0.212. This pattern extends to recall metrics, where reasoning LLMs show improvements of 0.243, 0.061, and 0.041 respectively, compared to non-reasoning LLMs’ more pronounced gains of 0.560, 0.345, and 0.135. We attribute this performance gap to the “overthinking” phenomenon (Cuadron et al., 2025; Sui et al., 2025), where excessive internal reasoning impedes effective action execution, a limitation that we examine in detail in the following subsection.
@@ 5.2 Tool Usage Analysis
Figure 2 presents the number of times each LLM invokes actions on the full dataset with Sci-Reproducer. We observe the following:
  • For code-related actions, reasoning LLMs demonstrate limited tool usage, employing “SearchFile”, “SearchCodeItem”, and “SearchWeb” only 25.0, 3.3, and 0.0 times on average, respectively. Non-reasoning LLMs use these same actions far more extensively, with averages of 210.4, 68.2, and 16.8 times respectively. This disparity reveals a fundamental behavioral difference: reasoning models favor internal deliberation over external information gathering. Conversely, reasoning LLMs invoke “Compiler” more frequently, indicating they require more debugging iterations due to inadequate contextual information gathering. This over-reliance on internal reasoning undermines performance: advanced models like o3-mini-high and o3-mini-low achieve execution accuracy comparable to GPT-4o-mini, negating their theoretical computational advantages.
  • Paper-related actions exhibit a similar pattern. Reasoning LLMs use “SearchPaper”, “SearchSection”, and “SearchLiterature” an average of 56.3, 70.0, and 20.0 times respectively, while non-reasoning LLMs demonstrate substantially higher usage at 244.8, 188.4, and 58.0 times respectively. Additionally, we observe a clear preference for target paper extraction over external literature consultation. Actions targeting the primary paper (“SearchPaper” and “SearchSection”) are invoked 174.1 and 144 times on average, significantly more than “SearchLiterature” which accesses related works only 43.8 times.
@@ 5.3.1 Syntax Errors
Table A4 shows the syntax error rates for each model across different configurations. Without the Code Agent, syntax errors occurred at rates of 80.3% (“NoAgent”) and 83.3% (“Paper Agent”). After implementing the Code Agent, these error rates dropped significantly to 29.4% (“Code Agent”) and 24.9% (“Sci-Reproducer”). The remaining syntax errors mainly result from incorrectly using repository dependencies. This occurs because our approach, unlike human developers, cannot dynamically access runtime information through a compiler during the code generation process.
@@ 6 Conclusion
We evaluate LLMs’ ability to replicate algorithms described in recent NLP papers. To support this, we introduce SciReplicate-Bench, a benchmark with rich annotations, and Sci-Reproducer, a dual-agent framework for bridging algorithm understanding and code generation. We assess performance using reasoning graph accuracy and standard implementation metrics. Results show the task is highly challenging, with failures largely caused by missing or inconsistent algorithm descriptions.
@@ 7 Acknowledgements
This work was supported in part by the UK Engineering and Physical Sciences Research Council (EPSRC) through a Turing AI Fellowship (grant no. EP/V020579/1, EP/V020579/2). We thank the authors of the selected papers for making their code openly available.
@@
Appendix A Details of the Annotation Process
Title Conference
1. From Zero to Hero: Cold-Start Anomaly Detection (Reiss et al., 2024) Findings of ACL 2024
2. Addressing Order Sensitivity of In-Context Demonstration Examples in Causal Language Models (Xiang et al., 2024) Findings of ACL 2024
3. Breaking the Ceiling of the LLM Community by Treating Token Generation as a Classification for Ensembling (Yu et al., 2024) Findings of EMNLP 2024
4. Simple but Effective Compound Geometric Operations for Temporal Knowledge Graph Completion (Ying et al., 2024) ACL 2024
5. When is Tree Search Useful for LLM Planning? It Depends on the Discriminator (Chen et al., 2024d) ACL 2024
6. Functional Overlap Reranking for Neural Code Generation (To et al., 2023) Findings of ACL 2024
7. Beyond Single-Event Extraction: Towards Efficient Document-Level Multi-Event Argument Extraction (Liu et al., 2024b) Findings of ACL 2024
8. Unifying Dual-Space Embedding for Entity Alignment via Contrastive Learning (Wang et al., 2024a) COLING 2025
9. Exploring Concept Depth: How Large Language Models Acquire Knowledge and Concepts at Different Layers? (Jin et al., 2024) COLING 2025
10. TRANSMI: A Framework to Create Strong Baselines from Multilingual Pretrained Language Models for Transliterated Data (Liu et al., 2024c) COLING 2025
11. Enhancing Knowledge Distillation of Large Language Models through Efficient Multi-Modal Distribution Alignment (Peng & Zhang, 2024) COLING 2025
12. Document-level Claim Extraction and Decontextualisation for Fact-Checking (Deng et al., 2024) ACL 2024
13. IRCAN: Mitigating Knowledge Conflicts in LLM Generation via Identifying and Reweighting Context-Aware Neurons (Shi et al., 2024) Neurips 2024
14. RouterDC: Query-Based Router by Dual Contrastive Learning for Assembling Large Language Models (Chen et al., 2024b) Neurips 2024
15. Unsupervised Homography Estimation on Multimodal Image Pair via Alternating Optimization (Song et al., 2024) Neurips 2024
16. RAPTOR: Recursive Abstractive Processing for Tree-Organized Retrieval (Sarthi et al., 2024) ICLR 2024
17. Less is KEN: a Universal and Simple Non-Parametric Pruning Algorithm for Large Language Models (Mastromattei & Zanzotto, 2024) Findings of ACL 2024
18. Adaptive Contrastive Search: Uncertainty-Guided Decoding for Open-Ended Text Generation (Arias et al., 2024) Findings of EMNLP 2024
19. MiniCheck: Efficient Fact-Checking of LLMs on Grounding Documents (Tang et al., 2024a) EMNLP 2024
20. Nearest Neighbor Normalization Improves Multimodal Retrieval (Chowdhury et al., 2024) EMNLP 2024
21. Neuron-Level Knowledge Attribution in Large Language Models (Yu & Ananiadou, 2023) EMNLP 2024
22. RaTEScore: A Metric for Radiology Report Generation (Zhao et al., 2024) EMNLP 2024
23. GREEN: Generative Radiology Report Evaluation and Error Notation (Ostmeier et al., 2024) Findings of EMNLP 2024
24. Style-Specific Neurons for Steering LLMs in Text Style Transfer (Lai et al., 2024) EMNLP 2024
25. Language-Specific Neurons: The Key to Multilingual Capabilities in Large Language Models (Tang et al., 2024b) ACL 2024
26. Reasoning Paths Optimization: Learning to Reason and Explore From Diverse Paths (Chia et al., 2024) Findings of EMNLP 2024
27. Lifelong Knowledge Editing for LLMs with Retrieval-Augmented Continuous Prompt Learning (Chen et al., 2024a) EMNLP 2024
28. NeuroMax: Enhancing Neural Topic Modeling via Maximizing Mutual Information and Group Topic Regularization (Pham et al., 2024) Findings of EMNLP 2024
29. Bridging Local Details and Global Context in Text-Attributed Graphs (Wang et al., 2024c) EMNLP 2024
30. Advancing Adversarial Suffix Transfer Learning on Aligned Large Language Models (Liu et al., 2024a) EMNLP 2024
31. SafeDecoding: Defending against Jailbreak Attacks via Safety-Aware Decoding (Xu et al., 2024) ACL 2024
32. MaskLID: Code-Switching Language Identification through Iterative Masking (Kargaran et al., 2024) ACL 2024
33. Towards Robust and Generalized Parameter-Efficient Fine-Tuning for Noisy Label Learning (Kim et al., 2024) ACL 2024
34. Learning to Maximize Mutual Information for Chain-of-Thought Distillation (Chen et al., 2024c) Findings of ACL 2024
35. GLiNER: Generalist Model for Named Entity Recognition using Bidirectional Transformer (Zaratiana et al., 2023) NAACL 2024
36. End-to-End Beam Retrieval for Multi-Hop Question Answering (Zhang et al., 2023a) NAACL 2024
Table A1: List of papers used in our Sci-Replicate benchmark.
@@ 21. Neuron-Level Knowledge Attribution in Large Language Models (Yu & Ananiadou, 2023) EMNLP 2024
We curated NLP papers from leading conferences in 2024, including ACL, EMNLP, ICLR, Neurips, and COLING. Using a web crawler, we collected accepted paper titles and employed the PapersWithCode API [^5] https://paperswithcode.com/api/v1/docs/ to identify those with open-source implementations. For each identified paper, we retrieved corresponding GitHub repository links and metadata (e.g., stars, issues, release dates) via the GitHub REST API [^6] https://docs.github.com/en/rest?apiVersion=2022-11-28.
@@
Finally, annotators created verification suites with 10 test cases per task, drawn from the original datasets used in each repository for the majority of papers. For a small number of repositories, fewer than 10 test cases could be constructed. For instance, algorithms that analyze LLM parameters may have only a single test case. Given the inherent randomness in many NLP implementations and potential machine-related variability, we addressed reproducibility from two angles:
@@
  • Controlling hardware variability: users were instructed to run both reference and generated code locally to eliminate discrepancies caused by hardware differences.
@@
Appendix B Details of the Task Categories
@@
Appendix C Details of the Actions
@@
For each Python file in the code repository, we utilize the Python AST  [^8] https://docs.python.org/3/library/ast.html package to parse the file and extract all defined classes, functions, and global variables. Unlike embedding-based code search methods (Zhang et al., 2024c; 2023c), the Code Agent in our framework directly provides the name of a code item. The tool then returns the corresponding definition if it exists; otherwise, it returns an empty response.
@@
Appendix D Human Evaluation of Reasoning Graph Accuracy
Table A2: Reasoning Graph Accuracy of Sci-Reproducer Models Evaluated by LLMs and Human Annotators
Sci-Reproducer Model Evaluation Method
GPT-4o Human (Mean ± Std)
GPT-4o-mini 0.782 0.768 (0.066)
O3-mini-medium 0.790 0.789 (0.033)
Table A3: Pearson Correlation between LLM and Human Judgments of Reasoning Graph Accuracy
Number of Data Points r P-value
20 0.7518 <<0.005
To validate the reliability of our LLM-based evaluation metrics, we conducted a human evaluation study on a subset of our benchmark. Three PhD students in computer science independently assessed the reasoning graph accuracy for 20 tasks generated by GPT-4o-mini and O3-mini-medium using Sci-Reproducer.
As shown in Table A2, human evaluations demonstrate strong alignment with our automated LLM-based assessments. The mean scores show consistent agreement between human annotators and GPT-4o evaluations, with standard deviations indicating reasonable inter-annotator consistency.
Furthermore, we computed the Pearson correlation coefficient between human and LLM-based evaluations across all assessed instances. As presented in Table A3, the correlation (r = 0.7518, p ¡ 0.005) indicates a strong positive relationship between human judgments and automated assessments, supporting the validity of our LLM-based evaluation approach.
While this preliminary validation demonstrates the effectiveness of our automated metrics, comprehensive human evaluation across the full benchmark remains an important direction for future work.
@@ Appendix E Figures and Tables
Figure A3: The overview of the SciReplicate-Bench.